OpenStudy (anonymous):

Solve the system of equations using substitution. x-8y=-58, 3y-5x=31

OpenStudy (anonymous):

x-8y=-58 x=8y-58 3y-5x=31 3y-5(8y-58)=31 3y-40y+290 -37y=-259 y=7 x-8y=-58 x-8(7)=-58 x-56=-58 x=-2 x=-2 y=7

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

No problem